OpenStudy (anonymous):

2a(7a2 – 3a + 9) Multiply and simplify? Help me and I'll give you a cookie... :3

OpenStudy (anonymous):

\[2a(7a^2 – 3a + 9)\] \[2a\times 7a^2-2a\times 3a+2a\times 9\] is a start

OpenStudy (anonymous):

then \[14a^3-6a^2+18a\] finishes the cookie

OpenStudy (anonymous):

Thanks c:

OpenStudy (anonymous):

yw